OpenStudy (anonymous):

How to make this problem so that all the #s have positive exponents : (6^-1)(r^-3)(r^2)/r^5

jimthompson5910 (jim_thompson5910):

Hint: use the idea that \[\LARGE x^{-k}=\frac{1}{x^{k}}\] where k is some positive number

OpenStudy (anonymous):

wat??????????

OpenStudy (anonymous):

x and k has nothing to do with it?!?!

OpenStudy (anonymous):

nvm its ok thanks anyways

OpenStudy (anonymous):

ill medal u

jimthompson5910 (jim_thompson5910):

For instance, if x = 6 and k = 1, then \[\LARGE x^{-k} = 6^{-1}\]

OpenStudy (anonymous):

ok kk

jimthompson5910 (jim_thompson5910):

So because \[\LARGE x^{-k}=\frac{1}{x^k}\], we can say that \[\LARGE 6^{-1}=\frac{1}{6^1}=\frac{1}{6}\] So \[\LARGE 6^{-1}=\frac{1}{6}\]

OpenStudy (anonymous):

ooohhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

thnx :)